Presented below are three independent situations.

(a) Flint Co. sold $2,020,000 of 12%, 10-year bonds at 104 on January 1, 2017. The bonds were dated January 1, 2017, and pay interest on July 1 and January 1. If Flint uses the straight-line method to amortize bond premium or discount, determine the amount of interest expense to be reported on July 1, 2017, and December 31, 2017. (Round answer to 0 decimal places, e.g. 38,548.)

Interest expense to be recorded $


(b) Buffalo Inc. issued $580,000 of 9%, 10-year bonds on June 30, 2017, for $480,209. This price provided a yield of 12% on the bonds. Interest is payable semiannually on December 31 and June 30. If Buffalo uses the effective-interest method, determine the amount of interest expense to record if financial statements are issued on October 31, 2017. (Round intermediate calculations to 6 decimal places, e.g. 1.251247 and final answer to 0 decimal places, e.g. 38,548.)

Interest expense to be recorded $
Answer #1

a) Premium on bonds payable = 2020000*4% = 80800

Interest expense = Interest paid-Premium amortized = (2020000*12%*6/12)-(80800/20) = $117160

Interest expense to be recorded = $117160

b) Interest expense on October 31,2017 = 480209*12%*4/12 = $19208

Interest expense to be recorded = $19208